OpenStudy (anonymous):

ln(x)+ln(x+2)=ln(8) i started with ln(x(x+2)=ln(8) so I have ln(x^2+2x)=ln(8) what next?

OpenStudy (lgbasallote):

subtract ln 8 from both sides ln (x^2 + 2x) - ln 8 = 0

OpenStudy (lgbasallote):

do you see the next step now?

OpenStudy (anonymous):

I tryed that then it would be ln (x^2+2x)/8 then how do i solve from there

OpenStudy (lgbasallote):

ln (x^2 +2x)/8 = 0 that = 0 is very important

OpenStudy (lgbasallote):

now you turn it into exponential form...ln has a base e right? so in exponential form it becomes \[e^0 = \frac{x^2+2x}{8}\] agree?

OpenStudy (anonymous):

ok i got that so e^0 equals 1 so then i can just solve for x?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

and it cannot be negative correct?

OpenStudy (lgbasallote):

yep..because you have an ln x

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

x+2 cannot be negative or equal to 0 as well so x has to be greater than 2